answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Fair space of two dice?

The sample space of 2 dice is 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, and 12.

In how many ways can you draw a club or face card from a deck of cards?

If the card is drawn at random, there are 25 ways (counting Aces as face cards).

What is events that have the same probability?

They are "events that have the same probability". Nothing more, nothing less.

When you mutilpy two numbers the outcome is?

When you multiply two numbers the outcome is called the product.

What symbol can't be repeated or subtracted?

V, L and D (5, 50 and 500 respectively). Repetition and subtraction are redundant because VV would be X, LL would be C, DD would be M, VX would be V, LC would be L and DM would be D. Although you could theoretically subtract V from L to get VL=45, the accepted convention does not permit any subtraction of 5, 50 or 500. 45 would be more correctly formed with XLV.

What are the odds of rolling 6 of any specific number using 6 dice in one roll?

The odds of rolling any specific number in one roll of one die is 1 in 6. Each die is unrelated, so the odds of rolling the same specific number using six dice in one roll is 1 in 6 to the 6th, or 1 in 46,656.

When rolling two dice what is the least likely number?

Assuming you're talking about standard six-sided dice numbered 1-6, the least likely rolls are 2 and 12, each of which should come up 1 time in every 36 rolls on average.

What is ACE Market?

This is a Stock Market in Kuala Lumpur Stock Exchance (KLSE) is known Bursa Malaysia. ACE is secondary stock market in KLSE. ACE's old name in KLSE is Mesdaq Market. ACE's mean; A is "access", C is "Certainty" and E is "Efficiency".

Taking over from the Mesdaq Market, the ACE Market is different because it will be sponsor-driven and open to companies of all sectors.

If a die is rolled 375 times and 5 occurs 80 times what is the probability of rolling a 5?

The probability of rolling a 5, based on the information given, is 80/375 or 16/75. Your problem describes a relative frequency approximation of probability.

When rolling a fair die 50 times what is the probability of rolling a 3 18 times?

Each roll of the die is a Bernoulli trial with a probability of success of 1/6 or about 0.1667. The probability of rolling a 3 exactly 18 times out of 50 is

18 ~ Bin(50, 1/6), which equals

(50!/(18!(50-18)!))(0.1667^18)(1-0.1667)^(50-18), which is about 0.00052.

Is a portion of the sample space?

Would you please be sure to complete your question when you resubmit it.

What does fifth - 9fifth percentile mean?

Suppose we were considering the heights of all 18-year-old women in a certain European city. We might put the heights in a (long) list and sort them from smallest to largest.

Let's suppose that there were exactly 500 women on the list. 5% of them would be 25. The height of the woman at position 25 from the top would be called the 25th percentile. 95% would be 475, and the height of the woman at position 475 would be called the 95th percentile.

Then we might say that the 5th to the 95th percentile would include all the women from position 25 through position 475.

What is the probability of flipping a coin a tail on the first and a head on the second or a head on the first and a tail on the second?

Two flips can have four possible results:

T - T

T - H

H - T

H - H

The question defines 'success' as either H-T or T-H.

Two successes out of four possibilities = probability of 2/4 = 1/2 = 50% .